OpenStudy (anonymous):

if Point B lies on AC. The length of AB is (3x - 2), the length of BC is (x - 7), and the length of AC is 15. Find the value of x.

OpenStudy (anonymous):

(3x-2)+(x-7)=15 4x-9=15 4x=24 x=6

OpenStudy (nurali):

AB + BC = AC (3x - 2)+(x - 7)=15 4x-9=15 4x=24 x=6